OpenStudy (anonymous):

A sample space consists of five simple events, E1,2,E3,E4,E5 -I figured out the first part of the question but the second part is where i'm stuck Part two: if P(E1)=3P(E2)=0.3, find the probability of the remaining simple events if you know that the remaining simple events are equally likely.

OpenStudy (kirbykirby):

\(P(E_1)=0.3\) \(2P(E_2)=0.3 \implies P(E_2)=\dfrac{0.3}{2}=0.15\) Now, the sum of all simple events of the sample space must sum to 1 to be a probability distribution. Hence, \(P(E_1)+P(E_2)+P(E_3)+P(E_4)+P(E_5)=1\) You already know from the question that \(P(E_3)=P(E_4)=P(E_5)\) since they are equally likely... you could essentially say that is the same as \(3P(E_3)\). Substitute that into the previous equation :)

OpenStudy (kirbykirby):

Ok I misread, it actually says \(3P(E_2)=0.3 \implies P(E_2)=\dfrac{0.3}{3}=0.1\)